OpenStudy (anonymous):

Here are my steps, could you tell me where I'm wrong, I know it is in my method of cross cancelling I just cant think of correct way: f(x)=−(3 5 )x+4 if f(3) −(3 5 )(3)+4 −(3(3) 5 )+4 −(9 5 )+4 −(9 5 )(5)+4(5) −9+20 11

OpenStudy (amistre64):

f(x)=−(3/ 5 )x+4 if f(3) −(3/ 5 )(3)+4 −(3(3)/ 5 )+4 −(9 /5 )+4 −(9 [5]? )(5)+4(5)/[5] <- right in here −(9+20)/5 11/5

OpenStudy (amistre64):

and keeping the "-" outside is bad karma

OpenStudy (amistre64):

\[f(x)=−\frac{3}{5}x+4 \ \ \ \ at\ f(3)\] \[=−\frac{3}{5}(3)+4 \] \[=\frac{−3(3)}{5}+4 \] \[=\frac{−9}{5}+4 \] \[=\frac{−9}{5}+4\frac{5}{5} \] \[=\frac{−9}{5}+\frac{20}{5} \] \[=\frac{−9+20}{5} \] \[=\frac{11}{5} \]

OpenStudy (anonymous):

Amazing, Thank you very much hopefully I can work through it with her now. Off to go see, wish me luck, lol

OpenStudy (amistre64):

good luck